# Yutong Xia[Uncaptioned image][Uncaptioned image]  Ao Qu[Uncaptioned image]  Yunhan Zheng[Uncaptioned image]  Yihong Tang[Uncaptioned image]  Dingyi Zhuang[Uncaptioned image] Yuxuan Liang[Uncaptioned image]  Shenhao Wang[Uncaptioned image]  Cathy Wu[Uncaptioned image]  Lijun Sun[Uncaptioned image] Roger Zimmermann[Uncaptioned image]  Jinhua Zhao[Uncaptioned image]

### v2

@@ Abstract
Urban causal research is essential for understanding the complex dynamics of cities and informing evidence-based policies. However, it is challenged by the inefficiency and bias of hypothesis generation, barriers to multimodal data complexity, and the methodological fragility of causal experimentation. Recent advances in large language models (LLMs) present an opportunity to rethink how urban causal analysis is conducted. This Perspective examines current urban causal research by analyzing taxonomies that categorize research topics, data sources, and methodological approaches to identify structural gaps. We then introduce an LLM-driven conceptual framework, AutoUrbanCI, composed of four distinct modular agents responsible for hypothesis generation, data engineering, experiment design and execution, and results interpretation with policy recommendations. We propose evaluation criteria for rigour and transparency, and reflect on implications for human–AI collaboration, equity, and accountability. We call for a new research agenda that embraces AI-augmented workflows not as replacements for human expertise, but as tools to broaden participation, improve reproducibility, and unlock more inclusive forms of urban causal reasoning. Equal contribution.
@@ 1 Introduction
Cities are complex systems at the heart of society’s most pressing challenges, from transportation congestion and housing affordability to climate resilience and equitable public services Ritchie et al. (2024); United Nations (2015); Xia et al. (2022). Scientific progress in urban research equips policymakers, planners, and communities with the knowledge to design evidence-based interventions that shape urban futures Batty (1997). Among these advances, causal inference lies at the heart of scientific urban research, offering tools to move beyond correlation and evaluates the true impacts of interventions Baum-Snow and Ferreira (2015). Causal inference methods have powered studies on how high-speed rail alters commuting patterns Heuermann and Schmieder (2019) how electric vehicle charging stations influence local economies Zheng et al. (2024a), and how vaccine mandates affect immunization uptake Karaivanov et al. (2022). Yet, despite its centrality, causal inference in urban contexts still faces structural limitations that restrict the scope, speed, and inclusivity of discovery.
@@
Current urban causal research faces three major challenges: 1) Hypothesis discovery is slow and biased: research questions are often manually crafted and shaped by disciplinary precedent or funding priorities, overlooking emerging issues and underrepresenting smaller cities or informal settlements Bell and Jayne (2009); Müller et al. (2025); 2) Data complexity impedes integration: while urban data is inherently multimodal (e.g., satellite imagery, mobility logs, public records), most studies rely on structured tabular data due to technical barriers, limiting the comprehensiveness of causal models Raghavan et al. (2020); Hong et al. (2023); 3) Experimentation is expensive and risk-prone: causal inference depends on stringent assumptions, such as valid counterfactuals or exogenous variation, and many studies are abandoned when these conditions fail, discouraging exploration of high-risk, high-reward policy questions Baum-Snow and Ferreira (2015); Matthay and Glymour (2022). The structural limitations of current urban causal research raise a timely question: Can artificial intelligence help expand the boundaries of urban scientific discovery?
@@
Recent advances in Large Language Models (LLMs) Achiam et al. (2023); Touvron et al. (2023); Team et al. (2023); Guo et al. (2025) and Multimodal Large Language Models (MLLMs) Wu et al. (2023) offer a compelling opportunity. LLMs demonstrate strong capabilities in understanding and generating text Radford et al. (2018), reasoning Wei et al. (2022); Jaech et al. (2024), planning Huang et al. (2022), code execution Chen et al. (2021), and tool use Schick et al. (2023); Qin et al. (2024). MLLMs extend these functions across diverse data modalities such as imagery Liu et al. (2023a), audio Chu et al. (2023), graph Tang et al. (2024), structured charts Han et al. (2023), and unstructured reports Li et al. (2024a), lowering the barriers to engaging with complex urban data at scale Liang et al. (2025). More recently, through automated hypothesis generation, integrated analysis of multimodal data, and adaptive causal study design, LLMs enable a shift from narrow, discipline-bound efforts to more inclusive and exploratory research Lu et al. (2024); Schmidgall et al. (2025); Gottweis et al. (2025).
@@
In this Perspective, we propose that LLMs can transform urban science by offer not just a means to simply assist or accelerat existing workflows, but a way to expand its epistemic boundaries. To explore this potential, we introduce AutoUrbanCI, an AI-powered framework for supporting causal inference in urban studies. AutoUrbanCI is a flexible, modular framework that supports the full urban causal pipeline—surfacing overlooked questions, assembling multimodal data, evaluating methodological assumptions, and generating interpretable outputs.
The outline of this paper are shown in Figure 1. Specifically, Section 2 presents a systematic review of urban causal research, identifying key gaps in data, methods, and geographic coverage. Section 3 introduces AutoUrbanCI, a modular agent-based framework that supports all stages of causal analysis, with a concrete multi-agent system use case. We also discuss potential challenges and propose evaluation methods. Section 4 reflects on how AutoUrbanCI enables broader participation from civic groups, governments, and urban communities, and offer a perspective on the human-AI collaboration paradigm. Finally, Section 5 summarizes the paper and outlines future directions. Our contributions are summarized as follows:
  • A Comprehensive Systematic Review. We conduct a large-scale empirical analysis of causal inference studies in urban research, covering over 2,400 articles across a decade from a flagship journal (i.e., Cities), identifying key trends and gaps in the use of causal inference.
  • A Modular Framework. We introduce AutoUrbanCI, a flexible, AI-augmented pipeline that supports all stages of causal analysis through modular agents spanning hypothesis generation, data preparation, experimental execution, and report writing, demonstrated through a concrete multi-agent system use case.
@@
  • Towards Inclusive Urban Science. We show how AutoUrbanCI reduces the entry barriers to rigorous urban causal analysis, allowing broader participation by civic organizations, municipal governments, and community groups - advancing the vision of a city “created by everybody” Jacobs (1961).
@@ 2 Landscape of Causal Inference in Urban Studies
This section provides a quantitative overview of the field based on a comprehensive empirical analysis. It identifies key trends in causal inference research across urban topics, data usage, methodological approaches, and geographic focus. These findings lay the foundation for designing AutoUrbanCI, our proposed AI framework for automated causal research, ensuring it is aligned with actual research needs, practices, and gaps (see Appendix C for methodological details).
@@ 2.4 Insights from Empirical Landscape
3 AutoUrbanCI framework
This section proposes AutoUrbanCI, a modular and flexible framework for urban causal inference. Specifically, we begin by identifying key barreiers of current manual workflows (Figure 4, left) and then highlight the capabilities of LLMs/MLLMs that make automation feasible (Figure 4, right). Building on these two foundations, we introduce a modular, agent-based framework that structures the causal inference process into four components (Figure 4, center), each supported by different operational modes and LLM capabilities. We then present a concrete multi-agent use case to illustrate the system in action. Finally, we discuss key challenges and outline an evaluation strategy for AutoUrbanCI.
@@ 3.1 Motivation
Figure 4: From limitations to design: motivating the AutoUrbanCI framework.. The challenges of manual urban causal workflows (left), the modular design of AutoUrbanCI offering agents with multiple operational modes for each stage (middle), and the capabilities of LLMs/MLLMs that can support distinct stages, denoted by colored checkmarks (right).
@@
Recent advances in large language and multimodal models (LLMs and MLLMs) have introduced new possibilities for automating causal inference workflows in urban research. We organize their relevant capabilities into two core categories: knowledge reasoning and procedural execution, which correspond closely to the intellectual and operational demands of the AutoUrbanCI framework.
@@
These capabilities provide the foundation for the agents within AutoUrbanCI, enabling both cognitive support (e.g., hypothesis refinement, narrative construction) and operational automation (e.g., data acquisition, model execution) across the entire urban causal inference pipeline, as illustrated in the right portion of Figure 4.
@@ 3.2 Framework structure and agent modes
Motivated by the above-mentioned limitations and emerging LLM/MLLMs’ capabilities, we propose AutoUrbanCI —a modular framework that structures urban causal inference into functional phases, with LLM agents assigned scoped roles and distinct modes of interaction. As shown in Figure 4 (middle), AutoUrbanCI is a flexible framework composed of four functional phases. Note that each phase can be supported by one or more agents, and agents can also operate across multiple phases as needed. Instead of a fixed pipeline, the framework provides a modular structure that adapts to varying data availability, task complexity, and automation needs. This flexibility allows AutoUrbanCI to support diverse use cases, including real-time policy decisions and data-scarce environments. The agent setup described here is one example; a concrete multi-agent instantiation is introduced in Section 3.3. We then describe each phase with different mode in detail below.
@@
The first phase of the causal workflow involves identifying meaningful and timely causal questions (i.e., hypothesis) rooted in urban phenomena. Traditionally, this step relies heavily on domain expertise, often guided by literature reviews or retrospective interpretations of policy outcomes—an approach that can introduce cognitive bias and limit scalability. To address this, AutoUrbanCI supports multiple operational modes, each tailored to different types of input sources for hypothesis generation.
@@ 3.3 Multi-agent system example
This section presents a concrete instantiation of AutoUrbanCI through a multi-agent system (MAS) example, illustrating how agents collaborate across the causal workflow in a realistic research scenario. We focus on a case study investigating the impact of congestion pricing on low-income commuters in New York City (NYC), as shown in Figure 5. This example demonstrates one possible path through the framework: Practice-Driven + Theory-Driven $\rightarrow$ Web-Based Retrieval $\rightarrow$ Quasi-Experimental Methods $\rightarrow$ Validity Check + Policy Narrative + Robustness Test. The system includes seven specialized agents—Urban Scientist, Reader, Data Engineer, Validator, Data Scientist, Experimenter, and Writer—each responsible for a distinct function in the pipeline. Detailed roles and responsibilities are described in Appendix D.
Figure 5: An example of the AutoUrbanCI system in practice. Phase 1 Hypothesis Generation: The Reader gathers policy news, social media posts, and academic papers, proposes a preliminary hypothesis, and refines it with the Urban Scientist while identifying the required data types. Phase 2 Dataset Preparation: The Data Engineer collects and preprocesses multimodal urban data. The Validator checks the validity of the resulting dataset. Phase 3 Experimental Design & Execution: The Experimenter runs exploratory regressions and conducts experiments under the design proposed by the Data Scientist. The Validator checks the robustness of the results. Phase 4 Results Interpretation: The Writer synthesizes results into a structured report in collaboration with the Data Scientist and Urban Scientist. The Validator checks the narrative alignment.
@@ 3.4 Challenges
We outline key challenges in realizing AutoUrbanCI as a scalable, policy-relevant system for causal analysis in urban environments.
@@
While AutoUrbanCI leverages LLM agents for flexible reasoning, these models are not inherently grounded in established causal methodologies. Without explicit constraints, agents may produce plausible-sounding but methodologically invalid conclusions, particularly in quasi-experimental designs that require strict adherence to identification assumptions. Moreover, imposing rigid workflow templates may limit adaptability to diverse research designs and non-standard causal frameworks Schmidgall et al. (2025). Effective alignment may involve expert-annotated prompting pipelines Wu et al. (2022) or fine-tuning on domain-specific corpora Liang et al. (2025). Ensuring that AutoUrbanCI agents reason within valid identification strategies, rather than relying on free-form generation, is essential for producing trustworthy and reproducible policy insights.
@@ 3.5 Evaluation
Evaluating the expert-level capabilities of LLM-driven systems is inherently challenging Si et al. (2024). The evaluation of AutoUrbanCI framework presents additional complexities, as it integrates multimodal data processing while maintaining the rigor required for causal inference experiments. Unlike general LLM-based automated research systems, which are primarily designed for machine learning tasks or scientific discovery in natural sciences, AutoUrbanCI introduces a novel paradigm tailored for urban causal research. This underscores the need to revisit the evaluation metrics for existing automated research systems, requiring evaluating their capacity to handle multimodal data, uphold the methodological rigor of causal inference, and ultimately generate meaningful causal insights that can inform urban policy and planning.
@@
For AutoUrbanCI, in addition to the existing evaluation methods, its inherent nature requires several additional evaluation focuses. We highlight three critical aspects: hypothesis generation significance, data processing quality, experiment design and causal inference validity. First, for idea generation evaluation, in addition to novelty and feasibility, it should also consider scientific rigor (i.e., ensuring asking appropriately clear and specific research questions adheres to established causal principles Bailey et al. (2024)) and interdisciplinary value (i.e., assessing whether they integrate insights from multiple domains, such as urban science, economics, and environmental studies, to provide innovative perspectives). Second, for data processing quality, its evaluation can focus on data completeness, consistency, preprocessing accuracy, scalability, and computational efficiency, ensuring robust, efficient, and effective multimodal data integration. Third, experiment design requires assessing the system’s ability to select appropriate causal identification strategies (e.g., DiD, IV), justify assumptions, and choose valid baselines. One practical evaluation approach is to reproduce classic urban causal inference studies and compare whether the agent’s design choices align with or diverge from established methodological best practices. Lastly, causal inference evaluation assesses causal structure feasibility, intervention validity, counterfactual consistency, and robustness to confounding. This can be achieved through synthetic data benchmarking, expert validation, or curated datasets derived from existing urban causal inference studies, ensuring that discovered causal insights are both statistically sound and practically relevant for urban policy applications. Table 2 summarizes the full set of evaluation dimensions and highlights the ones that are particularly specific to the AutoUrbanCI setting.
@@
Table 2: Evaluation dimensions for AutoUrbanCI. The final column marks criteria that are particularly specific to AutoUrbanCI.

### v3

@@ Abstract
Urban causal research is essential for understanding the complex, dynamic processes that shape cities and for informing evidence-based policies. However, current practices are often constrained by inefficient and biased hypothesis formulation, challenges in integrating multimodal data, and fragile experimental methodologies. Imagine a system that automatically estimates the causal impact of congestion pricing on commute times by income group or measures how new green spaces affect asthma rates across neighborhoods using satellite imagery and health reports, and then generates comprehensive, policy-ready outputs, including causal estimates, subgroup analyses, and actionable recommendations. In this Perspective, we propose UrbanCIA, an LLM-driven conceptual framework composed of four distinct modular agents responsible for hypothesis generation, data engineering, experiment design and execution, and results interpretation with policy insights. We begin by examining the current landscape of urban causal research through a structured taxonomy of research topics, data sources, and methodological approaches, revealing systemic limitations across the workflow. Next, we introduce the design principles and technological roadmap for the four modules in the proposed framework. We also propose evaluation criteria to assess the rigor and transparency of these AI-augmented processes. Finally, we reflect on the broader implications for human–AI collaboration, equity, and accountability. We call for a new research agenda that embraces LLM-driven tools as catalysts for more scalable, reproducible, and inclusive urban research. Equal contribution.
@@ 1 Introduction
Cities are complex systems shaped by interconnected forces, from transportation congestion and housing affordability to climate resilience and equitable public services Ritchie et al. (2024); United Nations (2015); Xia et al. (2022). Scientific progress in urban research equips policymakers, planners, and communities with the knowledge to design evidence-based interventions that shape urban futures Batty (1997). At the heart of this scientific effort is the pursuit of causal understanding: uncovering how policies, infrastructures, and behaviors interact to produce measurable impacts Baum-Snow and Ferreira (2015). Urban causal research typically follows a structured pipeline: beginning with a hypothesis, integrating diverse data sources, selecting appropriate methods, and interpreting results to inform decisions. For example, in assessing New York City’s congestion pricing policy News (2024), scientists may hypothesize that its impact on commute times differs by neighborhood income levels. This triggers a series of steps: collecting multimodal data such as subway logs, taxi trips, social media, and census records; applying methods like difference-in-differences with fixed effects Hernan and Robins (2020). The final analysis reveals not only whether commute times decreased, but whether the policy disproportionately benefits certain socioeconomic groups. Such approaches have also enabled studies on the effects of high-speed rail Heuermann and Schmieder (2019), electric vehicle infrastructure Zheng et al. (2024a), and vaccine mandates Karaivanov et al. (2022).
@@
However, this process remains labor-intensive, fragile, and unevenly accessible. We identify three major challenges: 1) Hypothesis discovery is slow and biased: research questions are often manually crafted and shaped by disciplinary precedent or funding priorities, overlooking emerging issues and underrepresenting smaller cities or informal settlements Bell and Jayne (2009); Müller et al. (2025); 2) Data complexity impedes integration: although urban environments generate rich and diverse data streams (e.g., satellite imagery, mobility logs, text reports), integrating them into causal studies remains technically demanding, leading many studies to rely primarily on structured tabular datasets Raghavan et al. (2020); Hong et al. (2023); 3) Experimentation is expensive and risk-prone: causal inference depends on stringent assumptions, such as valid counterfactuals or exogenous variation, and many studies are abandoned when these conditions fail, discouraging exploration of high-risk, high-reward policy questions Baum-Snow and Ferreira (2015); Matthay and Glymour (2022). Taken together, these challenges underscore the need for more systematic, data-aware, and technology-enabled approaches to causal inference in urban contexts. The structural limitations of current urban causal research raise a timely question: Can artificial intelligence help expand the scope, efficiency, and inclusivity of urban scientific discovery?
@@
Recent advances in Large Language Models (LLMs) Achiam et al. (2023); Touvron et al. (2023); Team et al. (2023); Guo et al. (2025) and Multimodal Large Language Models (MLLMs) Wu et al. (2023) offer new possibilities for reimagining how urban causal research is conducted. LLMs exhibit strong capabilities in language understanding Radford et al. (2018), reasoning Wei et al. (2022); Jaech et al. (2024), planning Huang et al. (2022), code execution Chen et al. (2021), and tool use Schick et al. (2023); Qin et al. (2024). MLLMs extend these functions across diverse data modalities such as imagery Liu et al. (2023a), audio Chu et al. (2023), graph Tang et al. (2024), structured charts Han et al. (2023), and unstructured reports Li et al. (2024a), lowering the barriers to engaging with complex urban data at scale Liang et al. (2025). Despite these technical advances, their application to urban causal research remains fragmented, often limited to isolated tasks such as data cleaning, code generation, or report drafting. Recent developments in agent-based LLM systems Lu et al. (2024); Schmidgall et al. (2025); Gottweis et al. (2025) suggest a promising direction: not using LLMs as end-to-end black boxes, but as specialized components within a coordinated reasoning system.
@@
In this Perspective, we argue that LLMs offer a path to reimagine how we ask questions, extract evidence, and generate causal knowledge in cities - at scale, and with broader reach. To explore this potential, we introduce UrbanCIA, a flexible and modular framework for Urban Causal Inference with Agents. UrbanCIA is a conceptual framework that supports the full urban causal pipeline: surfacing overlooked questions, assembling multimodal data, evaluating methodological assumptions, and generating interpretable outputs. Compared to conventional manual workflows and isolated AI-assisted tools, UrbanCIA offers a structured, extensible, and interpretable framework that supports the full urban causal inference process while preserving space for human judgment, domain expertise, and ethical deliberation.
The outline of this paper are shown in Figure 1. Specifically, Section 2 presents a systematic review of urban causal research, identifying key gaps in data, methods, and geographic coverage. Section 3 introduces UrbanCIA, a modular agent-based framework that supports all stages of causal inference, with a concrete multi-agent system use case. We also discuss potential challenges and propose evaluation methods. Section 4 reflects on how UrbanCIA enables broader participation from civic groups, governments, and urban communities, and offer a perspective on the human-AI collaboration paradigm. Finally, Section 5 summarizes the paper and outlines future directions. Our contributions are summarized as follows:
  • A Comprehensive Systematic Review. We conduct a large-scale empirical analysis of causal inference studies in urban research, covering over 2,400 articles across a decade from a flagship journal (i.e., Cities), identifying key trends and gaps in the use of causal inference under the urban context.
  • A Modular Framework. We introduce UrbanCIA, a flexible, AI-augmented pipeline that supports all stages of causal analysis through modular agents spanning hypothesis generation, data preparation, experimental execution, and report writing, demonstrated through a concrete multi-agent system use case.
@@
  • Towards Inclusive Urban Science. We show how UrbanCIA reduces the entry barriers to rigorous urban causal research, allowing broader participation by civic organizations, municipal governments, and community groups - advancing the vision of a city “created by everybody” Jacobs (1961).
@@ 2 Landscape of Causal Inference in Urban Studies
This section provides a quantitative overview of the field based on a comprehensive empirical analysis. It identifies key trends in causal inference research across urban topics, data usage, methodological approaches, and geographic focus. These findings lay the foundation for designing UrbanCIA, our proposed AI framework for automated causal research, ensuring it is aligned with actual research needs, practices, and gaps (see Appendix C for methodological details).
@@ 2.4 Insights from Empirical Landscape
3 UrbanCIA framework
This section proposes UrbanCIA, a modular and flexible framework for urban causal inference. Specifically, we begin by identifying key barreiers of current manual workflows (Figure 4, left) and then highlight the capabilities of LLMs/MLLMs that make automation feasible (Figure 4, right). Building on these two foundations, we introduce a modular, agent-based framework that structures the causal inference process into four components (Figure 4, center), each supported by different operational modes and LLM capabilities. We then present a concrete multi-agent use case to illustrate the system in action. Finally, we discuss key challenges and outline an evaluation strategy for UrbanCIA.
@@ 3.1 Motivation
Figure 4: From limitations to design: motivating the UrbanCIA framework.. The challenges of manual urban causal workflows (left), the modular design of UrbanCIA offering agents with multiple operational modes for each stage (middle), and the capabilities of LLMs/MLLMs that can support distinct stages, denoted by colored checkmarks (right).
@@
Recent advances in large language and multimodal models (LLMs and MLLMs) have introduced new possibilities for automating causal inference workflows in urban research. We organize their relevant capabilities into two core categories: knowledge reasoning and procedural execution, which correspond closely to the intellectual and operational demands of the UrbanCIA framework.
@@
These capabilities provide the foundation for the agents within UrbanCIA, enabling both cognitive support (e.g., hypothesis refinement, narrative construction) and operational automation (e.g., data acquisition, model execution) across the entire urban causal inference pipeline, as illustrated in the right portion of Figure 4.
@@ 3.2 Framework structure and agent modes
Motivated by the above-mentioned limitations and emerging LLM/MLLMs’ capabilities, we propose UrbanCIA —a modular framework that structures urban causal inference into functional phases, with LLM agents assigned scoped roles and distinct modes of interaction. As shown in Figure 4 (middle), UrbanCIA is a flexible framework composed of four functional phases. Note that each phase can be supported by one or more agents, and agents can also operate across multiple phases as needed. Instead of a fixed pipeline, the framework provides a modular structure that adapts to varying data availability, task complexity, and automation needs. This flexibility allows UrbanCIA to support diverse use cases, including real-time policy decisions and data-scarce environments. The agent setup described here is one example; a concrete multi-agent instantiation is introduced in Section 3.3. We then describe each phase with different mode in detail below.
@@
The first phase of the causal workflow involves identifying meaningful and timely causal questions (i.e., hypothesis) rooted in urban phenomena. Traditionally, this step relies heavily on domain expertise, often guided by literature reviews or retrospective interpretations of policy outcomes—an approach that can introduce cognitive bias and limit scalability. To address this, UrbanCIA supports multiple operational modes, each tailored to different types of input sources for hypothesis generation.
@@ 3.3 Multi-agent system example
This section presents a concrete instantiation of UrbanCIA through a multi-agent system (MAS) example, illustrating how agents collaborate across the causal workflow in a realistic research scenario. We focus on a case study investigating the impact of congestion pricing on low-income commuters in New York City (NYC), as shown in Figure 5. This example demonstrates one possible path through the framework: Practice-Driven + Theory-Driven $\rightarrow$ Web-Based Retrieval $\rightarrow$ Quasi-Experimental Methods $\rightarrow$ Validity Check + Policy Narrative + Robustness Test. The system includes seven specialized agents—Urban Scientist, Reader, Data Engineer, Validator, Data Scientist, Experimenter, and Writer—each responsible for a distinct function in the pipeline. Detailed roles and responsibilities are described in Appendix D.
Figure 5: An example of the UrbanCIA system in practice. Phase 1 Hypothesis Generation: The Reader gathers policy news, social media posts, and academic papers, proposes a preliminary hypothesis, and refines it with the Urban Scientist while identifying the required data types. Phase 2 Dataset Preparation: The Data Engineer collects and preprocesses multimodal urban data. The Validator checks the validity of the resulting dataset. Phase 3 Experimental Design & Execution: The Experimenter runs exploratory regressions and conducts experiments under the design proposed by the Data Scientist. The Validator checks the robustness of the results. Phase 4 Results Interpretation: The Writer synthesizes results into a structured report in collaboration with the Data Scientist and Urban Scientist. The Validator checks the narrative alignment.
@@ 3.4 Challenges
We outline key challenges in realizing UrbanCIA as a scalable, policy-relevant system for causal analysis in urban environments.
@@
While UrbanCIA leverages LLM agents for flexible reasoning, these models are not inherently grounded in established causal methodologies. Without explicit constraints, agents may produce plausible-sounding but methodologically invalid conclusions, particularly in quasi-experimental designs that require strict adherence to identification assumptions. Moreover, imposing rigid workflow templates may limit adaptability to diverse research designs and non-standard causal frameworks Schmidgall et al. (2025). Effective alignment may involve expert-annotated prompting pipelines Wu et al. (2022) or fine-tuning on domain-specific corpora Liang et al. (2025). Ensuring that UrbanCIA agents reason within valid identification strategies, rather than relying on free-form generation, is essential for producing trustworthy and reproducible policy insights.
@@ 3.5 Evaluation
Evaluating the expert-level capabilities of LLM-driven systems is inherently challenging Si et al. (2024). The evaluation of UrbanCIA framework presents additional complexities, as it integrates multimodal data processing while maintaining the rigor required for causal inference experiments. Unlike general LLM-based automated research systems, which are primarily designed for machine learning tasks or scientific discovery in natural sciences, UrbanCIA introduces a novel paradigm tailored for urban causal research. This underscores the need to revisit the evaluation metrics for existing automated research systems, requiring evaluating their capacity to handle multimodal data, uphold the methodological rigor of causal inference, and ultimately generate meaningful causal insights that can inform urban policy and planning.
@@
For UrbanCIA, in addition to the existing evaluation methods, its inherent nature requires several additional evaluation focuses. We highlight three critical aspects: hypothesis generation significance, data processing quality, experiment design and causal inference validity. First, for idea generation evaluation, in addition to novelty and feasibility, it should also consider scientific rigor (i.e., ensuring asking appropriately clear and specific research questions adheres to established causal principles Bailey et al. (2024)) and interdisciplinary value (i.e., assessing whether they integrate insights from multiple domains, such as urban science, economics, and environmental studies, to provide innovative perspectives). Second, for data processing quality, its evaluation can focus on data completeness, consistency, preprocessing accuracy, scalability, and computational efficiency, ensuring robust, efficient, and effective multimodal data integration. Third, experiment design requires assessing the system’s ability to select appropriate causal identification strategies (e.g., DiD, IV), justify assumptions, and choose valid baselines. One practical evaluation approach is to reproduce classic urban causal inference studies and compare whether the agent’s design choices align with or diverge from established methodological best practices. Lastly, causal inference evaluation assesses causal structure feasibility, intervention validity, counterfactual consistency, and robustness to confounding. This can be achieved through synthetic data benchmarking, expert validation, or curated datasets derived from existing urban causal inference studies, ensuring that discovered causal insights are both statistically sound and practically relevant for urban policy applications. Table 2 summarizes the full set of evaluation dimensions and highlights the ones that are particularly specific to the UrbanCIA setting.
@@
Table 2: Evaluation dimensions for UrbanCIA. The final column marks criteria that are particularly specific to UrbanCIA.